# Sensing Gravity with Polarized Electromagnetic Radiation

### v1

@@ Abstract
Polarization wiggling is a direct gravitational effect exerted by a gravitational field on any electromagnetic radiation traversing it. This effect is investigated in linear gravity for spacetimes with flat or conformally flat backgrounds. First, we show how the polarization wiggle rates can be computed in the conformal frame and transformed to the physical frame. Then it is shown that polarization wiggling is not sensitive to scalar perturbations to the metric, while vector and tensor perturbations do induce polarization wiggling. This poses two natural questions: Can polarized electromagnetic radiation be used to measure vectorial and tensorial components of gravitational fields directly? And if so, how? Next, polarization wiggling is studied for an arbitrary vector perturbation to the spacetime metric. In a stationary spacetime, the polarization wiggle rate is proportional to the difference in frame dragging rates between radiation emission and measurement events. We show how this can be used to measure the angular momentum of a gravitational source if the emitter orbits the gravitational source on a known orbit. Next, the effect of a gravitational tensor mode with arbitrary polarization is studied. Finally, the polarization wiggling effect induced by a gravitational tensor mode is analyzed. Two cases are demonstrated: A spacetime with a flat Minkowski background and an expanding cosmology with a conformally flat background. In both cases, the polarization wiggling frequency equals the frequency of the gravitational tensor mode, while the other state parameters are encoded in  the polarization wiggling amplitude and phase of the polarized radiation. We show that measurements of polarization wiggling frequency, amplitude and phase of polarized radiation from different directions enables all state parameters of a gravitational tensor mode to be determined.
@@ I Introduction
Electromagnetic radiation is our primary tool for probing the universe, and its polarization degrees of freedom represent a separate channel of information about the processes and phenomena with which the radiation interacts. While gravitational effects on electromagnetic radiation like gravitational lensing and redshift are well understood and used in astrophysical measurements, gravity’s direct effect on electromagnetic polarization is among the most elusive effects of gravity. Pending sufficiently accurate techniques for measuring gravitational effects on polarized radiation, it is important to understand them and how they can be used as separate means of probing gravitational fields, possibly beyond what is possible to do by other means.
@@
Defined in distinctly operational terms, polarizaton wiggling, on the other hand, provides an unambiguous quantification of the effect of gravity on the state of polarization of electromagnetic radiation [23]. It represents the effect from the perspective of a single inertial observer: Given a pointlike inertial emitter of polarized electromagnetic radation, it uniquely defines how the presence of a gravitational field along the trajectory of the radiation is manifested in observations made by this observer. Given an inertial emitter of polarized electromagnetic radiation, polarization wiggling manifests as a change in the measured state of polarization between subsequent observation events by a single inertial detector. For linearly or elliptically polarized radiation, it will be seen as a motion of the polarization axis. Polarizaton wiggling is by definition an observable gravitational effect on the state of electromagnetic polarization. It is induced by gravitational fields present along the radiation path between the radiation source and the detector.
In a recent paper, the polarization wiggling effect was studied for an arbitrary 4-dimensional spacetime and electromagnetic radiation with arbitrary state of polarization and frequency composition [23]. In the present paper, these results will be applied to study the effect of weak gravitational fields on electromagnetic polarization in more detail by using linearized gravity.
@@
In Section IV, we study polarization wiggling in linear gravity. In linear gravity, the gravitational field can be described as a perturbation to a homogeneous spacetime geometry referred to as the background geometry. When linearizing a metric theory of gravity, the gravitational field can be decomposed into scalar, vector and tensor perturbations that are sourced and evolve independently. In Section V, we show how this metric decomposition leads to a gauge invariant decomposition of the polarization wiggle rate into independent components induced by scalar, vector and tensor perturbations separately.
In Section VII, we consider the effect of general vector perturbations, and show how the polarization wiggle rate is related to the gravitomagnetic field. As a particular example, we calculate the effect of a stationary, axi-symmetric gravitational field induced by a rotating body on polarized radiation emitted by an emitter orbiting the gravitational source. We show how a series of polarization wiggle rate measurements of radiation emitted at different positions along the orbit will enable the angular momentum of the gravitational source to be determined.
@@ II Polarization Wiggling
In this section, we will briefly review the polarization wiggling effect. This effect was studied in detail in a recent paper [23], and we will use the nomenclature and notation introduced there. For simplicity, we will restrict our attention to plane electromagnetic waves with frequencies above the geometric optics limit. The results generalize to composite electromagnetic radiation with arbitrary state of polarization and frequency spectrum, given that the emitted frequency spectrum has negligible power below the geometric optics limit and that the radiation is emitted with a constant frequency spectrum [23].
Polarization wiggling is a motion (wiggling) of the polarization axis of electromagnetic radiation, as observed by an inertial observer. Polarization wiggling is a pure gravitational effect that is induced by the gravitational fields present along the geodesic segment between the emitter and the detector.
First, let us assume the presence of a ray of polarized electromagnetic radiation, represented as an electromagnetic plane wave. In the geometric optics limit, it can be represented by a vector potential
@@
This decomposition can be made for any spacetime geometry, and does not put any constraints on the spacetime geometry.
@@
The transverse polarization vector $\epsilon^{\mu}$ is defined as the transverse projection of the polarization vector $f^{\mu}$: $\epsilon^{\mu}\equiv S_{\nu}^{\mu}f^{\nu}$. Two quantities quantifying the gravitational effect on electromagnetic polarization are the polarization axis wiggle rate $\omega$ and the mean helicity wiggle rate $\chi$. They are defined as two independent projections of the transverse time rate of change of the transverse polarization vector $\epsilon^{\mu}$, and therefore quantify the time rate of change of the observed state of polarization in an inertial frame [23]. For linearly polarized radiation or radiation emitted with constant elliptical polarization, $\omega$ equals the rotation rate of the polarization axis. $\chi$ quantifies, as the name suggests, the  mean of the rates of change of the two helicity phases of the radiation. For radiation emitted with constant elliptical polarization, $\chi$ is proportional to the rotation rate of the polarization axis. $\omega$ is independent of both radiation frequency and the emitted state of polarization, while $\chi$ depends on the degree of circular polarization, $\mathcal{V}$. $\chi$ vanishes for linearly polarized radiation, while $\omega$ is undefined for radiation with pure circular polarization. Their transport equations along the null geodesic are
@@
A linear combination of eqs. (8) and (9) yields
@@ III The Scale Dependency of the Polarization Wiggle Rates
Here, $\tilde{g}$ denotes the metric of the physical spacetime geometry $\tilde{\mathcal{M}}$, while $g$ represents the metric of the conformal frame geometry $\mathcal{M}$. For example, the Minkowski geometry is the conformal frame of the conformally flat background geometries commonly used in cosmology.
We can make use of two facts to simplify the computation of polarization wiggle rates: 1) Null geodesics are invariant under conformal transformations, and 2) Maxwell’s equations in vacuum are conformally invariant. As will be shown shortly, this enables us to do all the necessary computations in the conformal frame, which is normally a much simpler geometry than the physical frame, and simply transform the results to the physical frame by using simple scale factors involving the conformal factor $\Omega$.
@@ III.2 Conformal Transformation of the Curvature Twist
Eq. (27) is remarkably simple. It complies with the intuitive notion that, if you scale a geometry by a fixed, coordinate-independent scale factor $\Omega$, the components of the Riemann tensor would be scaled by a factor $\Omega^{-2}$.
@@ III.3 The Scale Dependency of the Polarization Wiggle Rates
Eqs. (28) and (29) are generally valid for any spacetime with an arbitrary 4-dimensional conformal frame geometry and a homogeneous (time-dependent) conformal factor.
@@ V Scalar-Vector-Tensor Decomposition
$\omega^{\left(Q\right)}(\lambda)\equiv p\text{}\operatorname*{\int}\limits_{\lambda_{*}}^{\lambda}d\lambda^{\prime}\mathcal{Z}^{\left(Q\right)}(\lambda^{\prime})$ (41)
@@ V.1 Gauge Invariance of Curvature Twist
The curvature twist $\mathcal{Z}$ is by definition a scalar quantity, which means it is gauge invariant in linear gravity for an arbitrary metric perturbation $h_{\mu\nu}$. Since the metric perturbations $h_{\mu\nu}^{(S)},h_{\mu\nu}^{(V)}$ and $h_{\mu\nu}^{(T)}$ can be treated as independent perturbations, the corresponding curvature twist components $\mathcal{Z}^{(S)},\mathcal{Z}^{(V)}$ and $\mathcal{Z}^{(T)}$ are individially gauge invariant. This implies that the corresponding polarization wiggle rates $\omega^{(S)},\omega^{(V)}$ and $\omega^{(T)}$ are also gauge invariant. This is an important finding, because it means that we have found a way to physically quantify the effects of scalar, vector and tensor perturbations independently. Furthermore, we are free to choose metric gauge to evaluate the curvature twist and polarization wiggle rate components.
@@ V.2 Metric Decomposition
Let us do a scalar-vector-tensor decomposition of the metric perturbation $h_{\mu\nu}$into irreducible scalar vector and tensor components. We will apply transverse gauge. In this gauge the metric takes the form [24]
@@ VII.1 Polarization Wiggling from the Gravitomagnetic Field
In the transverse gauge, the vector component of the curvature twist is
@@
where $\mathcal{B}\equiv\hat{\text{\boldmath$p$}}\cdot\text{\boldmath$\mathcal{B}$}$. We will refer to $\mathcal{B}$ as the longitudinal gravitomagnetic field. The projection of the gravitomagnetic field $\mathcal{B}$ along the direction of propagation $\hat{p}$ induces a  rotation of the polarization axis of the radiation as it propagates along its null geodesic. This effect of the gravitomagnetic field on polarized electromagnetic radiation is referred to as the Gravitational Faraday Effect, due to its analogy with the Faraday Effect, which is the effect an ambient magnetic field has on the polarization of passing electromagnetic radiation. In other words, the Gravitational Faraday Effect can be viewed as polarization wiggling induced by vector perturbations through the resulting gravitomagnetic field.
@@
$\nu^{\left(u\right)}=-\frac{1}{2}\hat{\text{\boldmath$p$}}\cdot\text{\boldmath$\nabla$}\times\text{\boldmath$w$}=-\frac{1}{2}\mathcal{B}$ (45)
@@
$\omega^{\left(V\right)}(\lambda)=-\frac{1}{2}\left(p\operatorname*{\int}\limits_{\lambda_{*}}^{\lambda}d\lambda^{\prime}\frac{\partial\mathcal{B}}{\partial\eta}-\Delta\mathcal{B}\right),$ (46)
@@
We notice that, even if the gravitomagnetic field is stationary with $\partial\mathcal{B}/\partial\eta=0$, the vectorial contribution to the polarization axis wiggle rate, $\omega^{(V)}$, is still non-zero if the change in the longitudinal gravitomagnetic field between the emitter and detector is non-zero. The term $\Delta\mathcal{B}$ on the right-hand side of eq. (46) stems from the term $\nu^{(u)}$, and can be ascribed to the difference in inertial frame-dragging rate between the emitter and detector frames. This can be seen as follows: The angular velocity of the local inertial frame as measured by a distant observer in a rest frame is [28]
${\text{\boldmath{\hbox{\omega}}}}_{f}=-\frac{1}{2}\text{\boldmath{\hbox{\nabla}}}\times\text{\boldmath{\hbox{w}}}=-\frac{1}{2}\text{\boldmath{\hbox{\mathcal{B}}}}$
This  rotation of the local inertial frame is the manifestation of a gravitational effect known as the frame dragging effect, or dragging of inertial frames. For an emitter of radiation with polarization that is constant in the frame of emission, the polarization axis will rotate with the frame. The natural interpretation of eq. (46) is therefore that $\omega^{(V)}$ is the net angular speed of the polarization axis in the frame of observation after subtracting the difference $\Delta\mathcal{B}$ in angular frame dragging rates between the inertial frames of observation and emission.
@@
$\omega^{\left(V\right)}(\lambda)=-\frac{1}{2}\mathcal{B}(\lambda_{*})=\omega_{f}(\lambda_{*}),$ (47)
where $\omega_{f}\equiv\hat{\text{\boldmath$p$}}\cdot{\text{\boldmath$\omega$}}_{f}$ is the frame dragging rate around the direction of propagation. Conversely, given a distant emitter of polarized electromagnetic radiation with constant polarization at a position with no ambient gravitomagnetic field and a detector moving inertially in the vicinity of a gravitational source with a stationary gravitomagnetic field, the polarization axis wiggle rate of the measured polarized radiation is a direct measure of the angular frame dragging rate of the frame of observation:
$\omega^{\left(V\right)}(\lambda)=\frac{1}{2}\mathcal{B}(\lambda)=-\omega_{f}(\lambda).$ (48)
Finally, consider  eq. (46) for the case $\Delta\mathcal{B}=0$, which occurs when the gravitomagnetic field is negligible at both the detector and emitter locations. In this case, $\omega^{(V)}$ is non-zero only if the radiation traverses the vicinity of a gravitational source with a non-stationary gravitomagnetic field. The gravitomagnetic field is sourced by the matter current $j^{i}\equiv T^{0i}\simeq\rho v^{i}$ [29], where $T^{\mu\nu}$ is the energy-momentum tensor, $\rho$ is matter density and $v^{i}$ is matter velocity. Therefore, changing matter currents induce changes to the gravitomagnetic field, which in turn induce polarization wiggling in any passing electromagnetic radiation.
We conclude that the polarization axis wiggle rate $\omega^{(V)}$ is a direct measure of the longitudinal gravitomagnetic field $\mathcal{B}$. The contribution to the polarization axis wiggle rate from vector perturbations has two terms. The first contribution is due to the difference in the angular frame dragging rate between the frame of emission and the frame of observation. The second contribution comes from the presence of non-stationary gravitomagnetic fields along the radiation trajectory.
It is worth remarking that the polarization wiggle rate induced by a localized gravitomagnetic field on a Minkowski backgrond is to a large extent insensitive to distance between the emitter and detector. This can be seen from eq. (46). The second term on the right-hand side is clearly independent of distance, while the first term is independent of distance because $\mathcal{B}$ is non-zero in just a limited part of the path between emitter and detector. This characteristic distinguishes the polarization wiggle rate as a radiation observable from radiation intensity, which typically decreases as the radiation spreads with distance.

### v2

@@ Abstract
Polarization wiggling is an observational effect of a gravitational field on electromagnetic radiation traversing it. We show how polarization wiggle rates transform under homogeneous conformal transformations. This allows polarization wiggle rates to be computed in the conformal frame and transformed to the physical frame, which is useful when analyzing the effect in linear gravity for spacetimes with conformally flat backgrounds. In linear gravity, we find that the polarization wiggle rate contributions from scalar, vector and tensor perturbations are independent and gauge invariant. We show that polarization wiggling is not sensitive to scalar perturbations to the metric, while vector and tensor perturbations do induce polarization wiggling. This poses two natural questions: Can polarized electromagnetic radiation be used to measure vectorial and tensorial components of gravitational fields directly? And if so, how? Polarization wiggling is studied for an arbitrary vector perturbation to the spacetime metric. In a stationary spacetime, the polarization wiggle rate is proportional to the difference in frame dragging rate around the direction of propagation between radiation emission and measurement events. We show how this can be used to measure the angular momentum of a gravitational source if the emitter orbits the gravitational source on a known orbit. Finally, the polarization wiggling effect induced by a gravitational tensor mode with arbitrary polarization is analyzed. The effect is demonstrated for two specific cases: A spacetime with a flat Minkowski background and an expanding cosmology with a conformally flat background. In both cases, the polarization wiggling frequency equals the frequency of the gravitational tensor mode, while the other state parameters of the gravitational tensor mode are encoded in  the polarization wiggling amplitude and phase of the polarized radiation. We show that measurements of polarization wiggling frequency, amplitude and phase of polarized electromagnetic radiation emitted by sources at known positions in multiple directions enables all state parameters of a gravitational tensor mode to be determined.
@@ I Introduction
Electromagnetic radiation is our primary tool for probing the universe, and its polarization degrees of freedom represent a separate channel of information about the processes and phenomena with which the radiation interacts. While gravitational effects on electromagnetic radiation, like gravitational lensing and redshift, are well understood and used in astrophysical measurements, gravity’s direct effect on electromagnetic polarization is among the most elusive effects of gravity. Pending sufficiently accurate techniques for measuring gravitational effects on polarized radiation, it is important to understand them and how they can be used as separate means of probing gravitational fields, possibly beyond what is possible to do by other means.
@@
Polarizaton wiggling, on the other hand, provides an unambiguous quantification of the effect of gravity on the state of polarization of electromagnetic radiation because it is defined in distinctly operational terms [23]. It represents the effect from the perspective of a single inertial observer: Given a pointlike inertial emitter of polarized electromagnetic radation, it uniquely defines how the presence of a gravitational field along the trajectory of the radiation is manifested in observations made by this observer. Polarization wiggling manifests as a change in the measured state of polarization between subsequent observation events by a single inertial detector. For linearly or elliptically polarized radiation, it will be seen as a motion of the polarization axis in the frame of observation. Polarizaton wiggling is by definition an observable gravitational effect on the state of electromagnetic polarization. It is induced by gravitational fields present along the radiation path between the radiation source and the detector.
The polarization wiggling effect was described in a recent paper for a general 4-dimensional spacetime and electromagnetic radiation with an arbitrary state of polarization and frequency composition [23]. The present paper follows up by applying these results to study the effect of weak gravitational fields on electromagnetic polarization in more detail.
@@
In Section IV, we commence our study of polarization wiggling in linear gravity. In linear gravity, the gravitational field can be described as a perturbation to a homogeneous spacetime geometry referred to as the background geometry. When linearizing a metric theory of gravity, the gravitational field can be decomposed into scalar, vector and tensor perturbations that are sourced and evolve independently. In Section V, we show how this metric decomposition leads to a gauge invariant decomposition of the polarization wiggle rate into independent components induced by scalar, vector and tensor perturbations separately.
Section VI considers the effect of scalar perturbations to the spacetime metric. It is shown that scalar perturbations do not induce polarization wiggling.
In Section VII, we consider the effect of general vector perturbations, and show how the polarization wiggle rate is related to the gravitomagnetic field. As a particular example, we calculate the effect of a stationary, axi-symmetric gravitational field induced by a rotating body on polarized electromagnetic radiation emitted by an emitter orbiting the gravitational source. We show how a series of polarization wiggle rate measurements of radiation emitted at different positions along the orbit will enable the angular momentum of the gravitational source to be determined.
@@ II Polarization Wiggling
This section briefly reviews the polarization wiggling effect, described in a recent paper [23]. Both nomenclature and notation introduced there will be used in this paper. For simplicity, our attention will be restricted to plane electromagnetic waves with frequencies above the geometric optics limit. The results generalize to composite electromagnetic radiation with arbitrary state of polarization and frequency spectrum, given that the emitted frequency spectrum has negligible power below the geometric optics limit and that the radiation is emitted with a constant frequency spectrum [23].
Polarization wiggling is a motion (wiggling) of the polarization axis of electromagnetic radiation, as observed by an inertial observer. Polarization wiggling is a pure gravitational effect that is induced by the gravitational fields present along the null-geodesic segment between the emitter and the detector.
First, let us assume the presence of a ray of polarized electromagnetic radiation, represented as a plane electromagnetic wave. In the geometric optics limit, it can be represented by a vector potential
@@
The transverse polarization vector $\epsilon^{\mu}$ is defined as the transverse projection of the polarization vector $f^{\mu}$: $\epsilon^{\mu}\equiv S_{\nu}^{\mu}f^{\nu}$. Two quantities quantifying the gravitational effect on electromagnetic polarization are the polarization axis wiggle rate $\omega$ and the mean helicity wiggle rate $\chi$. They are defined as two transverse projections of the time rate of change of the transverse polarization vector $\epsilon^{\mu}$, and therefore quantify the time rate of change of the observed state of polarization in an inertial frame [23]
$\omega\equiv\mathrm{Re}\left\{\epsilon^{\gamma*}\epsilon_{\gamma\mu}\nabla_{u}\epsilon^{\mu}\right\}$
$\chi\equiv-i\epsilon^{\mu*}\nabla_{u}\epsilon_{\mu}.$
For linearly polarized radiation or radiation emitted with constant elliptical polarization, $\omega$ equals the rotation rate of the polarization axis. $\chi$ quantifies, as the name suggests, the  mean of the rates of change of the two helicity phases of the radiation. For radiation emitted with constant elliptical polarization, $\chi$ is proportional to the rotation rate of the polarization axis. $\omega$ is independent of both radiation frequency and the emitted state of polarization, while $\chi$ depends on the degree of circular polarization, $\mathcal{V}$. $\chi$ vanishes for linearly polarized radiation, while $\omega$ is undefined for radiation with pure circular polarization. Their transport equations along the null geodesic are
@@
By the use of eq. (10), a linear combination of eqs. (8) and (9) yields
@@ III The Scale Dependency of the Polarization Wiggle Rates
Here, $\tilde{g}$ denotes the metric of the physical spacetime geometry $\tilde{\mathcal{M}}$, while $g$ represents the metric of the conformal frame geometry $\mathcal{M}$. For example, Minkowski space is the conformal frame of the conformally flat background geometries commonly used in cosmology.
We can make use of two facts to simplify the computation of polarization wiggle rates: 1) Null geodesics are invariant under conformal transformations, and 2) Maxwell’s equations in vacuum are conformally invariant. As will be shown shortly, this enables us to, with a given  solution to the Einstein equations in the physical frame, do all the necessary computations in the conformal frame and transform the results to the physical frame. Since the conformal frame is typically a simpler geometry than the physical frame, the ability to compute wiggle rates in the conformal frame offers a significant simplification of these computations for spacetimes with non-trivial conformal factors, such as expanding cosmologies.
@@ III.2 Conformal Transformation of the Curvature Twist
Eq. (27) is remarkably simple. It complies with the intuitive notion that, if you scale a geometry by a constant, coordinate-independent scale factor $\Omega$, the components of the Riemann tensor will be scaled by a factor $\Omega^{-2}$.
@@ III.3 The Scale Dependency of the Polarization Wiggle Rates
Eqs. (28) and (29) are generally valid for any spacetime with an arbitrary 4-dimensional conformal frame geometry and a homogeneous (time-dependent) conformal factor. It is a remarkably simple result that enables us to compute the wiggle rates in the presumably much simpler conformal frame and transform them to the physical frame.
@@ V Scalar-Vector-Tensor Decomposition
$\omega^{\left(Q\right)}(\lambda)=p\text{}\operatorname*{\int}\limits_{\lambda_{*}}^{\lambda}d\lambda^{\prime}\mathcal{Z}^{\left(Q\right)}(\lambda^{\prime})$ (41)
@@ V.1 Gauge Invariance of Curvature Twist
The curvature twist $\mathcal{Z}$ is by definition a scalar quantity, which means it is gauge invariant in linear gravity for an arbitrary metric perturbation $h_{\mu\nu}$. Since the metric perturbations $h_{\mu\nu}^{(S)},h_{\mu\nu}^{(V)}$ and $h_{\mu\nu}^{(T)}$ can be treated as independent perturbations, the corresponding curvature twist components $\mathcal{Z}^{(S)},\mathcal{Z}^{(V)}$ and $\mathcal{Z}^{(T)}$ are individially gauge invariant. This implies that the corresponding polarization wiggle rates $\omega^{(S)},\omega^{(V)}$ and $\omega^{(T)}$ are also gauge invariant. This is an important finding, because it means that we have found a way to physically quantify the effects of scalar, vector and tensor perturbations on electromagnetic polarization independently. Furthermore, we are free to choose metric gauge to evaluate the curvature twist and polarization wiggle rate components.
@@ V.2 Metric Decomposition
Let us do a scalar-vector-tensor decomposition of the metric perturbation $h_{\mu\nu}$ in the transverse gauge. In this gauge, the metric takes the form [24]
@@ VII.1 Polarization Wiggling from the Gravitomagnetic Field
In the transverse gauge, the vector component of the curvature twist can be written
@@
where $\mathcal{B}\equiv\hat{\text{\boldmath$p$}}\cdot\text{\boldmath$\mathcal{B}$}$. We will refer to $\mathcal{B}$ as the longitudinal gravitomagnetic field. The projection of the gravitomagnetic field $\mathcal{B}$ along the direction of propagation $\hat{p}$ induces a  rotation of the polarization axis of the radiation as it propagates along its null geodesic. This effect of the gravitomagnetic field on polarized electromagnetic radiation is known as the Gravitational Faraday Effect, due to its analogy with the Faraday Effect, which is the effect an ambient magnetic field has on the polarization of passing electromagnetic radiation. In other words, the Gravitational Faraday Effect can be viewed as polarization wiggling induced by vector perturbations through the resulting gravitomagnetic field.
@@
$\nu^{\left(u\right)}=-\frac{1}{2}\hat{\text{\boldmath$p$}}\cdot\text{\boldmath$\nabla$}\times\text{\boldmath$w$}=-\frac{1}{2}\mathcal{B}.$ (45)
@@
$\omega^{\left(V\right)}(\lambda)=\frac{1}{2}\Delta\mathcal{B}-\frac{1}{2}p\operatorname*{\int}\limits_{\lambda_{*}}^{\lambda}d\lambda^{\prime}\frac{\partial\mathcal{B}}{\partial\eta},$ (46)
@@
We notice that, even if the gravitomagnetic field is stationary with $\partial\mathcal{B}/\partial\eta=0$, the vectorial contribution to the polarization axis wiggle rate, $\omega^{(V)}$, is still non-zero if the change in the longitudinal gravitomagnetic field between the emitter and detector is non-zero. The $\Delta\mathcal{B}$ term on the right-hand side of eq. (46) stems from the term $\nu^{(u)}$, and can be ascribed to the difference in inertial frame-dragging rate between the emitter and detector frames. This can be seen as follows: The angular velocity of the local inertial frame as measured by a distant observer in a rest frame is [28]
${\text{\boldmath{\hbox{\omega}}}}_{f}=-\frac{1}{2}\text{\boldmath{\hbox{\nabla}}}\times\text{\boldmath{\hbox{w}}}=-\frac{1}{2}\mathcal{B}.$ (47)
This  rotation of the local inertial frame is the manifestation of a gravitational effect known as the frame dragging effect, or dragging of inertial frames. For an emitter of radiation with polarization that is constant in the frame of emission, the polarization axis will rotate with the frame of emission and opposite of the frame of observation:
$\omega^{\left(V\right)}(\lambda)=\omega_{f}(\lambda_{*})+\Delta\omega-\omega_{f}(\lambda),$ (48)
where $\omega_{f}\equiv\hat{\text{\boldmath$p$}}\cdot{\text{\boldmath$\omega$}}_{f}$ is the frame dragging rate around the direction of propagation and
$\Delta\omega\equiv p\operatorname*{\int}\limits_{\lambda_{*}}^{\lambda}d\lambda^{\prime}\frac{\partial{\omega}_{f}(\lambda^{\prime})}{\partial\eta}$
is the net gain in angular rotation rate of an axis subject to an angular acceleration $\partial\omega_{f}/\partial\eta$ along the radiation trajectory. It is evident from eq. (47) that $\Delta\omega$ is non-zero only for non-stationary gravitomagnetic fields. The natural interpretation of eq. (48) is that $\omega^{(V)}$ is the net angular rotation rate of the polarization axis after subtracting the angular frame dragging rate of the frame of observation.
@@
$\omega^{\left(V\right)}(\lambda)=\omega_{f}(\lambda_{*}),$ (49)
Conversely, given a distant emitter of polarized electromagnetic radiation with constant polarization at a position with no ambient gravitomagnetic field and a detector moving inertially in the vicinity of a gravitational source with a stationary gravitomagnetic field, the polarization axis wiggle rate of the measured polarized radiation is a direct measure of the angular frame dragging rate of the frame of observation:
$\omega^{\left(V\right)}(\lambda)=-\omega_{f}(\lambda).$ (50)
Finally, consider  eq. (48) for the case $\omega_{f}(\lambda_{*})=\omega_{f}(\lambda)=0$, which occurs when the gravitomagnetic field is negligible at both the detector and emitter locations. In this case, $\omega^{(V)}$ is non-zero only if the radiation traverses the vicinity of a gravitational source with a non-stationary gravitomagnetic field. The gravitomagnetic field is sourced by the matter current $j^{i}\equiv T^{0i}\simeq\rho v^{i}$ [29], where $T^{\mu\nu}$ is the energy-momentum tensor, $\rho$ is matter density and $v^{i}$ is matter velocity. Therefore, changing matter currents induce changes to the gravitomagnetic field, which in turn induce polarization wiggling in any passing electromagnetic radiation.
We conclude that the polarization axis wiggle rate $\omega^{(V)}$ is a direct measure of the longitudinal gravitomagnetic field $\mathcal{B}$. The contribution to the polarization axis wiggle rate from vector perturbations has two terms. The first contribution is due to the difference in the angular frame dragging rate between the frame of emission and the frame of observation. The second contribution stems from any non-stationary gravitomagnetic fields present along the radiation trajectory.
It is worth remarking that the polarization wiggle rate induced by a localized gravitomagnetic field on a Minkowski backgrond is to a large extent insensitive to distance between the emitter and detector. This can be seen from eq. (46). The first term on the right-hand side is clearly independent of distance, while the second term is independent of distance because $\mathcal{B}$ is non-zero in just a limited part of the path between emitter and detector.